OpenStudy (anonymous):

The formula for electrical power (P, in watts) is p=V^2/R where V is the voltage in volts and R is the resistance in ohms. For a particular brand of water heater, R is 32 ohms and P is 0.5 watts. The formula to find V is V = PR V = (PR)^(1/2) V = P^2R , and the voltage at which the heater operates is 16 5 4 volts.

OpenStudy (anonymous):

V=PR V=(PR)^(1/2) V=P^2R

OpenStudy (irishboy123):